# Large-scale peculiar velocities in the universe

### v2

@@ 6.3.1 Two alternative perspectives
Let us consider a family of realistic observers moving relative to the reference (CMB) frame of the universe. Assuming that and are the 4-velocities associated with the former and the latter group of observers respectively, we have
In the above, is the peculiar velocity of the tilted (the realistic) observers with respect to the CMB frame, whereas is the velocity of the latter relative to the former (with and by construction – see Fig. 3). Also, when the peculiar motion is not relativistic (as it happens in our case), we have and . When studying the effects of relative motion, one is free to choose any of the above approaches. Here, we will adopt the perspective of (6.3.1a), but it goes without saying that nothing would have changed if we were to adopt that of (6.3.1b).
We also remind the reader that all the variables describing perturbations, including the peculiar-velocity field (), vanish identically in the FRW background. This satisfies the Stewart & Walker criterion for gauge invariance [215] and guarantees that our analysis is independent of the gauge-choice “connecting” the background and the perturbed spacetime.
@@ 6.3.2 The key role of the peculiar flux
In what follows, we will demonstrate the implications of the above arguments in perturbed FRW universes, starting with the Einstein-de Sitter background and then generalising to Friedmann models with nonzero spatial curvature. The latter studies will also allow us to probe the linear evolution of peculiar velocities in locally underdense and/or overdense regions of a nearly flat FRW universe.
@@
with zero linear pressure in both coordinate systems (i.e. – see (5.3.4b) in § 5.3).999Alternatively, one could set in the reference (CMB) frame of the idealised observers. In that case the linear expressions (5.3.4c) and (5.3.1b) respectively lead to and in the tilted frame of the matter. In other words, it is now the realistic observers that “see” the cosmic medium as imperfect and also “experience” a nonzero 4-acceleration (entirely due to relative-motion effects as well). Despite the change of perspective, the results remain unchanged, since it is only the relative motion between the two frames that matters (e.g. see [43, 44, 45] for the comparison). It is important to note that, according to Eq. (6.3.2a), the cosmic medium cannot be treated as perfect when peculiar motions are present. The “imperfection” appears as a nonzero peculiar flux vector (), solely triggered by the moving matter (e.g. see § 5.2.1 in [89] for related comments and also for the generalised linear version of (6.3.2a)).
In a proper general-relativistic study, the peculiar flux contributes to the energy-momentum tensor and, through Einstein’s equations, its gravitational input feeds into the conservation laws and eventually emerges in a host of relativistic relations. Among them, are the formulae governing the linear evolution of inhomogeneities in the density and in the expansion. These are respectively monitored by the spatial gradients and , which in the presence of peculiar motions evolve as
@@
respectively.101010There are more than one ways of obtaining (6.3.3) and (6.3.4). The faster is by linearising the nonlinear formulae (2.3.1) and (2.3.2) of [88], or equivalently expressions (10.101) and (10.102) of [89], while using the linear relations (6.3.2a) and (6.3.2b) as well. Note that, without the peculiar-velocity terms seen on the right-hand side of the above differential equations, expressions (6.3.3) and (6.3.4) coincide with those used in studies free of peculiar-motion effects [88, 89]. These velocity terms reflect the general-relativistic input of the peculiar flux to the gravitational field (see § 6.3.2 before).
@@
6.3.3 Peculiar velocities in FRW universes with
@@ 6.3.4 The role of the peculiar 4-acceleration
6.3.5 Peculiar velocities in FRW universes with
Having tested the case of the Einstein-de Sitter background, let us move on to Friedmann universes with non-Euclidean spatial geometry. Allowing for nonzero spatial curvature in the background, makes it possible to study the evolution of peculiar velocities in the underdense and overdense environments of the open and the closed FRW universes respectively. The downside is that the complexity of the analytic treatments increases considerably. To begin with, in Friedmann universes with nonzero spatial curvature, it is necessary to express the evolution equations in terms of the conformal time (, with ), instead of proper time (e.g. see [208] for further discussion). This is achieved by means of the transformation formulae
@@ 6.3.5 Peculiar velocities in FRW universes with
with at all times. On the other hand, the scale factor of a closed FRW universe with dust is given by (see solution (4.3.14) in § 4.3.4) and the rescaled Hubble parameter is , with . In this environment, Eq. (6.3.10) reads
@@
The case of an open FRW background: The hyperbolic geometry of the open FRW background makes it essentially impossible to obtain analytic linear solutions for the whole evolution of these models. The existing analytic solutions usually cover the late-time evolution, when , and . At this limit, Eq. (6.3.11) simplifies to
@@
with still given by (6.3.18). The growing mode of the above corresponds to , which suggests that the lower the value of , the slower the linear growth of the peculiar-velocity field. It is also straightforward to show that as long as , while keeping in mind that at late times. Recall that the relativistic analysis has led to on spatially flat FRW backgrounds (see § 6.3.3 here).131313At late times, the open FRW model approaches the empty Milne universe. Then, taking the limit of (6.3.18), we obtain and , leading to the solution for the linear evolution of the peculiar-velocity field. Therefore, near the Milne limit, linear peculiar-velocities remain essentially constant, which agrees with solution (6.3.19) at the limit.
Although the above solutions cover limiting periods in the evolution of Friedmann universes with hyperbolic spatial geometry, they do suggest that linear peculiar velocities grow slower on open FRW backgrounds than on their spatially flat counterparts. In fact, the lower the background density parameter, the slower the resulting peculiar growth-rate.
The case of a closed FRW background: In contrast to their flat and open counterparts, Friedmann models with spherical spatial geometry are finite both in space and in time (conformal as well as proper). Another key difference is that the second term on the left-hand side of (6.3.12) changes sign when the closed FRW background enters its contracting phase. In particular, the coefficient of the aforementioned term is negative throughout the expanding phase (i.e. when ), but turns positive once the universe has started to collapse (i.e. for ). We therefore expect the peculiar velocities to evolve differently during these two phases. Next, we will put this expectation to the test by looking for analytic solutions at characteristic moments in the lifetime of a spatially closed Friedmann universe.
@@
guaranteeing one growing and one decaying mode (with and respectively). The growing mode ensures that the higher the value of the density parameter, the stronger the linear growth of the -field. Although (6.3.21) and (6.3.21) are formally identical to the late-time solution obtained in open FRW universes, there is a key difference between the two results, because in the case and here. This allows for growth rates stronger than those obtained in flat and open FRW models. In fact, when , we find that . The latter argues for an arbitrarily strong linear growth of the peculiar velocity, if the background density parameter is allowed to increase arbitrarily as well.
@@
As in the expanding phase before, and mark the growing and decaying modes respectively. Also, the larger/smaller the value of , the stronger/weaker the growth of the peculiar velocities. However, comparing the sets (6.3.21), (6.3.22) and (6.3.24), (6.3.25) closer, shows that peculiar-velocity perturbations grow faster during the expanding phase than when the universe is contracting. The reason is that the density parameter of the closed Friedmann models increases when they expand, but decreases when they contract (recall that for dust). This means that, unlike solution (6.3.21)-(6.3.22), the growing mode of (6.3.24)-(6.3.25) has a finite upper bound. Indeed, in a highly dense FRW background with , we find near the middle of its contraction.
At the moment of maximum expansion , and . Close to this “turning” point, Eq. (6.3.12) gives , suggesting linear growth (in terms of conformal time) for the -field. This solution is independent of the value of and also holds when the closed FRW background is replaced by the static Einstein model.141414On the Einstein-static background, the homogeneous part of Eq. (6.3.5) reduces to , which leads to on all scales (recall that in the Einstein-static universe). This confirms the instability of the Einstein universe against linear peculiar-velocity perturbations as well [46].151515The Einstein static universe has long been known to be unstable against linear homogeneous and isotropic density perturbations as well [216], though its instability is less straightforward in the case of conformal metric and inhomogeneous distortions [217, 218, 219]
@@
In summary, linear peculiar velocities can grow faster in closed FRW universes than in their flat counterparts and even faster than in the spatially open ones [46]. This is still a tentative claim, however, since our analytic solutions do not cover the whole lifetime of Friedmann models with . Having said that, there is a persistent pattern in all the solutions, which supports the aforementioned claim. Also note that, even if the universe is close to the Euclidean limit globally, our results suggest faster peculiar velocities in locally overdense regions (e.g. in galaxy clusters) than in underdense domains (e.g. in the so-called voids) [46].
@@ 6.5 Peculiar velocities and dark-energy
In accordance with the current cosmological model, namely with the CDM paradigm, the Einstein-de Sitter period of the post-recombination universe is followed by a late phase of accelerated expansion. The latter is triggered by the dominance of an as yet unknown and elusive matter field, commonly referred to as dark energy.
A proper analysis of the peculiar-velocity evolution during the recent dark-energy epoch requires a two-fluid approach, analogous to that given in § 8.2 below. More specifically, the material component of the late-time universe should be treated a mixture of the dominant dark-energy component and of the subdominant baryonic matter and/or CDM. Although, the detailed analysis is still missing, following [45] and the results presented in § 6.4.3 here, one would expect to see the linear peculiar velocities to grow at a slower pace relative to the preceding Einstein-de Sitter epoch, if not to decay.
When combining this expectation with the faster growing rates for the linear peculiar-velocity field of the relativistic analysis (see § 6.3 before), one expects to measure bulk-flow velocities faster than expected, but also to see them drop at relatively low redshifts, as the accelerated expansion starts to dominate the kinematics of our cosmos. Interestingly, bulk-flow profiles with a late-time velocity drop have been reported in [29] and more recently in [33] (see Fig. 12 and Fig. 13 in § 9.3 below). It is therefore conceivable that the bulk-velocity decay reported in [29, 33] could reflect/mark the onset of the recent dark-energy era. This possibility may be also supported by the fact that, so far at least, the surveys that report slower bulk velocities are confined to relatively low redshifts (corresponding to scales smaller than Mpc), whereas those reporting faster bulk flows extend beyond the aforementioned threshold (see related discussion in § 9 below).
@@ 7.1 Relative-motion effects on the deceleration parameter
Therefore, the difference between the deceleration parameter () measured locally by the bulk-flow observer and the one () measured in the CMB frame, namely the deceleration parameter of the universe as a whole, is determined by the time derivative of the local volume scalar (). The latter is given by the “peculiar” Raychaudhuri equation, which in a tilted almost-FRW universe takes the form [86, 212]
@@ 7.1.2 The Einstein-de Sitter case
After equipartition the universe is believed to be close to the Einstein-de Sitter model, which is a flat FRW cosmology with zero pressure. Perturbing the above spacetime and allowing for non-relativistic peculiar velocities, the associated 4-acceleration is given by the linear expression (7.1.10) above. Substituting the latter into the right-hand side of (7.1.7) gives

### v3

@@ 6.3.1 Two alternative perspectives
Let us consider a family of realistic observers moving relative to the reference (CMB) frame during the Enstein-de Sitter epoch of the universe. Assuming that and are the 4-velocities associated with the former and the latter group of observers respectively, we have
In the above, is the peculiar velocity of the tilted (the realistic) observers with respect to the CMB frame, whereas is the velocity of the latter relative to the former (with and by construction – see Fig. 3). We remind the reader that both the CMB and the “tilted” frames are located at the same spacetime event by construction (e.g. see [227] and § 5.1 here). This guarantees that a Lorentz boosting between these two coordinate systems is local and does not violate causality. Instead, the transformation relates the quantities measured by the realistic (tilted) observers to those measured by their idealised (CMB) parters. This in turn allows the realistic observers to quantify the kinematic “contamination” induced by their own peculiar motion. Also, when the peculiar motion is not relativistic (as it happens in our case), we have and . When studying the effects of relative motion, one is free to choose any of the above approaches. Here, we will adopt the perspective of (6.3.1a), but it goes without saying that nothing would have changed if we were to adopt that of (6.3.1b).
We also remind the reader that all the variables describing perturbations, including the peculiar-velocity field (), vanish identically in the FLRW background. This satisfies the Stewart & Walker criterion for gauge invariance [231] and guarantees that our analysis is independent of the gauge-choice “connecting” the background and the perturbed spacetime.
@@ 6.3.2 The key role of the peculiar flux
In what follows, we will demonstrate the implications of the above arguments in perturbed FLRW universes, starting with the Einstein-de Sitter background and then generalising to Friedmann models with nonzero spatial curvature. The latter studies will also allow us to probe the linear evolution of peculiar velocities in locally underdense and/or overdense regions of a nearly flat FLRW universe.
@@
with zero linear pressure in both coordinate systems (i.e. – see (5.3.4b) in § 5.3). It is important to note that, according to Eq. (6.3.2a), the cosmic medium cannot be treated as perfect when peculiar motions are present. The “imperfection” appears as a nonzero peculiar flux vector (), solely triggered by the moving matter (e.g. see § 5.2.1 in [100] for related comments and also for the generalised linear version of (6.3.2a)). Overall, while the pressureless dust follows a geodesic frame, where the 4-acceleration () vanishes, there is nonzero 4-acceleration () in the CMB frame, which acts as the driving force of the linear peculiar-velocity field (see Eq. (6.3.2b) above).
At this point, before proceeding any further, it is important to note that the linear expression (6.3.2b) is also the staring point of the quasi-Newtonian analysis (see Eq. (6.2.1) in § 6.2.2 previously). However, the agreement between the quasi-Newtonian and the relativistic approaches to the question of the peculiar-velocity evolution stops here.
In a proper relativistic study no zero-vorticity and zero-shear constraints are imposed upon the perturbed spacetime. Also, there is no need to express the 4-acceleration as the gradient of an ad hoc scalar potential ( - see expression (6.2.2) in § 6.2.2) and no need to use the ansatz to determine its temporal evolution. Most importantly, the gravitational input of the peculiar flux is fully accounted for throughout the analysis.
Via Einstein’s equations, the flux contribution to relativistic gravitational field feeds into the conservation laws and eventually emerges in a host of relativistic relations (see [51] for a recent streamlined presentation). Among them are the formulae governing the linear evolution of inhomogeneities in the density and in the expansion. These are respectively monitored by the spatial gradients and , which in the presence of peculiar motions evolve as
@@
respectively [51].101010In [51] Eqs. (6.3.3) and (6.3.4) were obtained by taking the spatial gradients of the energy-density conservation law and of the Raychaudhuri equation (see (4.2.9) and (4.2.3) respectively), both linearised in the CMB frame. There are more ways of obtaining (6.3.3) and (6.3.4) however. The faster is by linearising the nonlinear formulae (2.3.1) and (2.3.2) of [99], or equivalently expressions (10.101) and (10.102) of [100], while using the linear relations (6.3.2a) and (6.3.2b) as well. Note that, without the peculiar-velocity terms seen on the right-hand side of the above differential equations, expressions (6.3.3) and (6.3.4) coincide with those used in studies free of peculiar-motion effects [99, 100]. These velocity terms reflect the general-relativistic input of the peculiar flux to the gravitational field (see § 6.3.2 before).
@@
6.3.3 The relativistic 4-acceleration
Before proceeding to the solutions of (6.3.5), it is worth deriving the linear form of the relativistic 4-acceleration in the presence of peculiar motions and in the absence of matter pressure. In such an environment, the relativistic energy conservation law (see Eq. (4.2.9) in § 4.2.2 previously for the nonlinear expression) linearises to
Note the extra flux-related term on the right-hand side of the above, which results form the purely general relativistic contribution of the peculiar flux () to the gravitational field (see § 6.3.2 before). As we will show next, accounting for the implications of the aforementioned flux-term is what separates the relativistic studies of peculiar motions from the rest.
Taking the spatial gradient of (6.3.6), employing the linear commutation law , recalling that and keeping up to first-order terms, one arrives at the following linear expression for the 4-acceleration by [50]
The profound difference between the above and its quasi-Newtonian analogue (compare to Eq. (6.2.2) in § 6.2.2 before), predisposes for an analogous difference between the results of the two approaches (see § 6.3.4 next).
We would also like to draw the reader’s attention to the density and the expansion gradients seen on the right-hand side of (6.3.7) and represented by and respectively. The presence of these inhomogeneity gradients, in connection with Eq. (6.3.2b), ensures that linear peculiar velocities are driven by structure formation. Finally, combining the above with Eq. (6.3.2b) leads to expression (6.3.3).
6.3.4 Peculiar velocities in FLRW universes with
@@ 6.3.5 The role of the peculiar 4-acceleration
6.3.6 Peculiar velocities in FLRW universes with
Having tested the case of the Einstein-de Sitter background, let us move on to Friedmann universes with non-Euclidean spatial geometry. Allowing for nonzero spatial curvature in the background, makes it possible to study the evolution of peculiar velocities in the underdense and overdense environments of the open and the closed FLRW universes respectively. The downside is that the complexity of the analytic treatments increases considerably. To begin with, in Friedmann universes with nonzero spatial curvature, it is necessary to express the evolution equations in terms of the conformal time (, with ), instead of proper time (e.g. see [220] for further discussion). This is achieved by means of the transformation formulae
@@ 6.3.6 Peculiar velocities in FLRW universes with
with at all times. On the other hand, the scale factor of a closed FLRW universe with dust is given by (see solution (4.3.14) in § 4.3.4) and the rescaled Hubble parameter is , with . In this environment, Eq. (6.3.12) reads
@@
The case of an open FLRW background: The hyperbolic geometry of the open FLRW background makes it essentially impossible to obtain analytic linear solutions for the whole evolution of these models. The existing analytic solutions usually cover the late-time evolution, when , and . At this limit, Eq. (6.3.13) simplifies to
@@
with still given by (6.3.20). The growing mode of the above corresponds to , which suggests that the lower the value of , the slower the linear growth of the peculiar-velocity field. It is also straightforward to show that as long as , while keeping in mind that at late times. Recall that the relativistic analysis has led to on spatially flat FLRW backgrounds (see § 6.3.4 here).131313At late times, the open FLRW model approaches the empty Milne universe. Then, taking the limit of (6.3.20), we obtain and , leading to the solution for the linear evolution of the peculiar-velocity field. Therefore, near the Milne limit, linear peculiar-velocities remain essentially constant, which agrees with solution (6.3.21) at the limit.
Although the above solutions cover limiting periods in the evolution of Friedmann universes with hyperbolic spatial geometry, they do suggest that linear peculiar velocities grow slower on open FLRW backgrounds than on their spatially flat counterparts. In fact, the lower the background density parameter, the slower the resulting peculiar growth-rate.
The case of a closed FLRW background: In contrast to their flat and open counterparts, Friedmann models with spherical spatial geometry are finite both in space and in time (conformal as well as proper). Another key difference is that the second term on the left-hand side of (6.3.14) changes sign when the closed FLRW background enters its contracting phase. In particular, the coefficient of the aforementioned term is negative throughout the expanding phase (i.e. when ), but turns positive once the universe has started to collapse (i.e. for ). We therefore expect the peculiar velocities to evolve differently during these two phases. Next, we will put this expectation to the test by looking for analytic solutions at characteristic moments in the lifetime of a spatially closed Friedmann universe.
@@
guaranteeing one growing and one decaying mode (with and respectively). The growing mode ensures that the higher the value of the density parameter, the stronger the linear growth of the -field. Although (6.3.23) and (6.3.23) are formally identical to the late-time solution obtained in open FLRW universes, there is a key difference between the two results, because in the case and here. This allows for growth rates stronger than those obtained in flat and open FLRW models. In fact, when , we find that . The latter argues for an arbitrarily strong linear growth of the peculiar velocity, if the background density parameter is allowed to increase arbitrarily as well.
@@
As in the expanding phase before, and mark the growing and decaying modes respectively. Also, the larger/smaller the value of , the stronger/weaker the growth of the peculiar velocities. However, comparing the sets (6.3.23), (6.3.24) and (6.3.26), (6.3.27) closer, shows that peculiar-velocity perturbations grow faster during the expanding phase than when the universe is contracting. The reason is that the density parameter of the closed Friedmann models increases when they expand, but decreases when they contract (recall that for dust). This means that, unlike solution (6.3.23)-(6.3.24), the growing mode of (6.3.26)-(6.3.27) has a finite upper bound. Indeed, in a highly dense FLRW background with , we find near the middle of its contraction.
At the moment of maximum expansion , and . Close to this “turning” point, Eq. (6.3.14) gives , suggesting linear growth (in terms of conformal time) for the -field. This solution is independent of the value of and also holds when the closed FLRW background is replaced by the static Einstein model.141414On the Einstein-static background, the homogeneous part of Eq. (6.3.5) reduces to , which leads to on all scales (recall that in the Einstein-static universe). This confirms the instability of the Einstein universe against linear peculiar-velocity perturbations as well [50].151515The Einstein static universe has long been known to be unstable against linear homogeneous and isotropic density perturbations as well [232], though its instability is less straightforward in the case of conformal metric and inhomogeneous distortions [233, 234, 235]
@@
In summary, linear peculiar velocities can grow faster in closed FLRW universes than in their flat counterparts and even faster than in the spatially open ones [50]. This is still a tentative claim, however, since our analytic solutions do not cover the whole lifetime of Friedmann models with . Having said that, there is a persistent pattern in all the solutions, which supports the aforementioned claim. Also note that, even if the universe is close to the Euclidean limit globally, our results suggest faster peculiar velocities in locally overdense regions (e.g. in galaxy clusters) than in underdense domains (e.g. in the so-called voids) [50].
@@ 6.5 Peculiar velocities and dark-energy
In accordance with the current cosmological model, namely with the CDM paradigm, the Einstein-de Sitter period of the post-recombination universe is followed by a late phase of accelerated expansion. The latter is triggered by the dominance of an as yet unknown and elusive matter field, commonly referred to as dark energy. Intuitively speaking, once the -dominated phase of the universe starts, one expects the accelerated expansion to suppress the growth of the peculiar-velocity perturbation, as it does with essentially all types of distortions.
A proper analysis of the peculiar-velocity evolution during the recent dark-energy epoch requires a two-fluid approach, analogous to that given in § 8.2 below. More specifically, the material component of the late-time universe should be treated a mixture of the dominant dark-energy component and of the subdominant baryonic matter and/or CDM. Although, the detailed analysis is still missing, the numerical study of [253] reached the conclusion that dark energy inhibits the growth of peculiar velocities, when compared to the preceding Einstein-de Sitter era. This result is also corroborated by the theoretical analysis of [49], as outlined in § 6.4.3 here, which found exponential decay for the linear peculiar-velocity field during a period of de Sitter inflation. On these grounds, one would expect to see the linear peculiar velocities to grow at a slower pace relative to the Einstein-de Sitter epoch, if not to decay.
When combining this expectation with the faster growing rates for the linear peculiar-velocity field of the relativistic analysis (see § 6.3 before), one expects to measure bulk-flow velocities faster than anticipated, but also to see them drop at relatively low redshifts, as the accelerated expansion starts to dominate the kinematics of our cosmos [51]. Interestingly, bulk-flow profiles with a late-time velocity drop have been reported in [29] and more recently in [33] (see Fig. 1 in [51], as well as Fig. 13 and Fig. 14 in § 9.3 below). Although more work is required, it is conceivable that the bulk-velocity decay detected in [29, 33] could reflect/mark the onset of the recent dark-energy era. This possibility may be also supported by the fact that, so far at least, the surveys that report slower bulk velocities seem confined to relatively low redshifts (corresponding to scales smaller than Mpc), whereas those reporting faster bulk flows appear to extend beyond the aforementioned threshold (see related discussion in § 9 below).
We will close this section with a brief reference to the numerical treatments that aim to simulate the evolution structure formation. In the recent literature one can find a number of such works that use relativistic techniques in their study. Nevertheless, to the best of our knowledge, the available simulations do not focus on the evolution and the implications of the peculiar velocities reviewed here. Perhaps the only exception are a couple of works that simulated the rotational component of the peculiar-velocity field (e.g. see [254, 255]). However, the peculiar vorticity, as well as the peculiar shear, grow slower that the peculiar velocity itself [47], so their effects are considerably weaker. Moreover, almost all of the simulations operate during the recent -dominated phase of universe, where the accelerated expansion inhibits the growth of the peculiar-velocity field [254, 255, 256, 257, 258, 259]. Having said that, the advent of the relativistic numerical simulations is a very recent one and the codes are still under development. So, there is considerable room for improvement. In this respect, the exact analytical results reviewed in this section can provide a valuable testing ground for the existing and the upcoming simulations. The fact that the aforementioned analytical studies are also gauge invariant is an additional advantage, since their solutions are free of spurious gauge-depended modes. Recall that the numerical studies rely on particular gauge-choices to address the problems in hand.
@@ 7.1 Relative-motion effects on the deceleration parameter
Therefore, the difference between the deceleration parameter () measured locally by the bulk-flow observer and the one () measured in the CMB frame, namely the deceleration parameter of the universe as a whole, is determined by the time derivative of the local volume scalar (). The latter is given by the “peculiar” Raychaudhuri equation, which in a tilted almost-FLRW universe takes the form [97, 228]
@@ 7.1.2 The Einstein-de Sitter case
After equipartition the universe is believed to be close to the Einstein-de Sitter model, which is a flat FLRW cosmology with zero pressure. Perturbing the above spacetime and allowing for non-relativistic peculiar velocities, the associated 4-acceleration is given by the linear expression (7.1.10) above. Substituting the latter into the right-hand side of (7.1.7) gives